my room is 12 ft x 18 ft what is the area and perimeter?
Answer by matthew_sessoms(39)

First, you must know the formulas for "perimeter" and "area".
Perimeter (P) = 2l + 2w, as l=length and w=width.
Area (A) = lw


It doesn't really matter what measurement is length or width in this particular question. So I'll just let l=12 and w=18.


P = 2(12) + 2(18)
P = 60


A = (12)(18)
A = 216


So the answer is:
Perimeter = 60 ft.
Area = 216 ft^2
MS